8. Show that the pressure due to 760 mm of mercury (Hg) is equal to 101kN/m². The specific gravity of Hg is 13.6. (4)10 Determine the absolute pressure at a depth of 20 ft below the free surface of a body of water (Assume density of water is
Answer #1

1 ) Mercury head ginenth= 760mm = 0.76m Specific grauity of Hg = 13.6. Density (8)= 13,600 kg/m3. - granity g=10m/s Heres conzoft density of water = 62.4 16m/f43 Pabs-Patm = gauge v Pabs= Patm + Pgange. p=ggh ginen (39) = 62.41bm/ 543 - P= (62.4)(20)